what is the mass in grams of NH4Cl present in 50.0ml of a 0.800 M NH4Cl solution?
1
Expert's answer
2020-01-30T04:31:34-0500

Answer on Question #102006 – Chemistry – Other


Task:

What is the mass in grams of NH4Cl present in 50.0ml of a 0.800 M NH4Cl solution?


Solution:

1) Molarity (M) is the concentration of a solution expressed as the number of moles of solute per liter of solution:

Molarity (M) = moles solute / liters solution = CM = n / V

Then,

CM(NH4Cl) = n(NH4Cl) / V(NH4Cl);

n(NH4Cl) = CM(NH4Cl) * V(NH4Cl) = 0.800 M * 0.050 L = 0.04 moles.

2) Molar mass of NH4Cl:

M(NH4Cl) = Ar(N) + 4*Ar(H) + Ar(Cl) = 14 +4*1 + 35.5 = 53.5 (g/mol).

3) The relationship between molar mass, mass and moles can be expressed as a mathematical equation as shown below:

n = m / M

Then,

n(NH4Cl) = m(NH4Cl) / M(NH4Cl);

m(NH4Cl) = n(NH4Cl) * M(NH4Cl) = 0.04 mol * 53.5 g/mol = 2.14 g

m(NH4Cl) = 2.140 g.


Answer: 2.140 g is the mass in grams of NH4Cl.